OpenStudy (anonymous):

Rewrite the following equation in slope-intercept form: 2x – y = 6 a y= 2x + 6 b y = -2x – 6 c y = 2x + 6 d y = 2x – 6

OpenStudy (anonymous):

Remember, slope-intercept form is (y-k)=m(x-h). All you need to do is move terms around. Start by adding y to both sides:\[2x=y+6\] Then subtract 6 from both sides:\[2x-6=y \] or \[y=2x-6\]

OpenStudy (anonymous):

You can also go further and factor that answer so it will look like \[y=2(x-3)\]